[figure: 9.7 cm]

Effect of hydrogen on the magnetic characteristics of Nd_{2}Fe_{14}B single crystal

N.Yu. Pankratov^{a,c}, I.S. Tereshina^{a,c}, S.A.Nikitin^{a,c}, T. Palewski^{a}, H. Drulis^{b}, Yu.G. Pastushenkov^{d}, M.V.Makarova^{c} aInternational Laboratory for High Magnetic Fields and Low Temperatures, 95 Gajowicka str., 53-421 Wroclaw, Poland

bPolish Akademy of Sciences, W. Trzebiatowski Institute of Low Temperature and

Structure Research, P.O. Box 1410, 50-950 Wroclaw 2, Poland cM.V.Lomonosov Moscow State University, 119992 Moscow, Russia

dFaculty of Physics, Tver State University, 33 Zheljabova str., 170002 Tver, Russia

Nd_{2}Fe_{14}B single crystals were hydrogenated to the composition Nd_{2}Fe_{14}BH_{x} where x ranges from 0 to 3.3. Magnetic characteristics of the present single crystals were investigated in the temperature range 4.2 to 750 K. It is established that hydrogen absorption does not change the crystal structure of these compounds but is accompanied by an expansion of a unit cell volume. The measurements show that the hydrogen absorption leads to an increase of the Curie temperatures and values of the iron moment in all the hydrides investigated. A systematic study of the influence of hydrogen on the spin reorientation transition is reported. The part of the work performed by N.Yu. Pankratov was supported by INTAS fellowship grant N YSF 2001/2-92.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : N.Yu. Pankratov

Address for correspondence : M.V.Lomonosov Moscow State University, Department of Physics, KOFEF Vorobyevy Gory, 119992 Moscow, Russia

Email address : nik@rem.phys.msu.su